# Instruction-Following Pruning for Large Language Models

### v2

@@ 4.2 Evaluation with Input-Specific Masks
Second, IFPruning demonstrates superior performance compared to the structured pruning method. Please note that the structured pruning model, Pruning+Distill, benefits from additional training signals due to knowledge distillation from a larger model as the teacher. Nevertheless, IFPruning consistently outperforms this baseline. Our method achieves higher performance across a variety of benchmarks, including AlpacaEval, Arena Hard, math problems, coding tasks, tool use, MMLU, and most Core Text tasks.
Third, we observe a clear improvement in performance with the IFPruning method as the size of the LLM increases. As the source model size scales from 6B to 9B and then to 12B, there is a noticeable performance boost on most of the datasets.
@@
Table 2: Performance comparison between IFPruning with per-input masks and per-task masks.
We highlight the following findings. First, in the lower layers, especially the first layer, the LLM tends to activate very similar sub-networks for different inputs. As we move to higher layers, the parameter selection becomes more diverse. Second, as shown in the figure, IFPruning activates distinct sub-networks for different domains in higher layers. For instance, models pruned for MMLU-CS have substantial overlap with those for Code-Alpaca, significantly more than with other domains. Similarly, models for MMLU-Math, MMLU-Physics, and GSM8K share a high proportion of activated parameters, which diverge notably from the activation patterns for MMLU-History, MMLU-Law, and GPTTeacher. Third, the overlap rate along the diagonal of the heatmap is very high, reflecting the strong similarity between models for inputs within the same domain. Finally, as expected, GPTTeacher is an instruction-following dataset that covers a wide range of domains, therefore it does not exhibit significant domain-specific characteristics, resulting in a self-overlap rate that is lower than in other domains. In summary, the activated parameters are interpretable and reveal clear patterns in different domains. This interpretability aligns with the high performance of our method, as the LLM model dynamically activates the parameters most suitable for each input.
@@ 4.3 Evaluation with Task-Specific Masks
An additional noteworthy capability of IFPruning is task-specific pruning. While IFPruning can effectively prune a model on a per-input basis, the algorithm also demonstrates the ability to select one sub-network that can be used for different inputs within the same task or domain. Specifically, we evaluate our models on math problems, coding tasks, the MMLU dataset, and machine translation tasks from Flores-101 (Goyal et al., 2022). The examples from the same task or domains share the same instructions when selecting sub-networks. For MMLU, we manually merge similar subsets into one domain. For example, “astronomy”, “college physics”, “conceptual physics”, and “high school physics” subsets are merged as the “MMLU-physics” domain for evaluation. We list the subsets included in each MMLU domain in Appendix A.2.
@@
The performance is shown in Table 2. The task instructions (i.e., inputs to the sparsity predictor) for each dataset are provided in Table  4 in Appendix A.3. We compare the task-specific pruning capabilities of our algorithm with two methods: the dense LLM with 3B parameters and the standard IFPruning with per-input mask. Our key findings are as follows. First, IFPruning can generate high-quality task-specific masks without additional training. we observe substantial performance improvements of the task-specific IFPruning over the dense baseline across all datasets. When applying per-task masks, IFPruning still outperforms the dense LLM by 5%-12% on coding tasks, 5%-6% on MATH, and 1% - 5% on MMLU subsets. For translation tasks, IFPruning selects the appropriate sub-networks without additional training, achieving an improvement of 4.8 points in BLEU score on average. Note that these task-specific masks are directly predicted by the sparsity predictor, requiring no additional fine-tuning of either the sparsity predictor or the LLM for each task. This highlights the “zero-shot” pruning capability of our method. Second, compared to the standard input-specific IFPruning, the task-specific IFPruning exhibit minimal performance degradation across math problems, coding tasks, and MMLU. The performance gaps are mostly under 1%, indicating the robustness of our method.
@@ 5 Conclusion
6 Limitations
This paper still has several limitations that need to be addressed in future work. First, we focus on exploring the potential of dynamically pruning LLMs based on contextual information. While this approach is well suited for models on consumer-facing devices such as laptops and phones, additional challenges remain for server-side serving when the input is a batch of user requests containing different tasks. One possible solution is to cluster user requests so that requests within the same batch share similar activated sub-networks. Another possibility is learning to prune for a composition of multiple tasks. These are interesting questions for future work.
Second, the current method relies on end-to-end optimization. While this allows for simple and scalable training, it may not fully utilize the training examples. The performance could be improved by adopting more advanced training strategies. For instance, using contrastive loss could encourage higher overlap rates among sub-networks for similar inputs, making the model more robust.
7 Societal Impact
@@
Appendix A Example Appendix
A.1 Model Architecture
We list the detailed model architecture of the pre-trained models used in our experiments in Table 3. All models use the same model dimension, attention dimensions and the number of Transformer layers. The only difference is the feed-forward dimension. Accordingly, our method learns to select 6656 FFN dimensions from the 6B and 9B model.
@@ A.1 Model Architecture
A.2 Subsets included in MMLU domains
@@
A.3 Detailed Per-task Prompts
@@
Multiple-E Swift You are an expert Swift programmer. You will be given a Swift function definition in documentation comments after ///. Write the code to complete the function. Here is an example input: ```swift /// Write a swiftthon function to count inversions in an array. func get_Inv_Count(arr: [Int]) -> Int
HumanEval-Python You are an expert Python programmer. You will be given a Python function definition and some test examples in triple quotes """. Write the code which should pass the tests. Here is an example input: ```python def greatest_common_divisor(a: int, b: int) -> int: """Return a greatest common divisor of two integers a and b $>$ greatest_common_divisor(3, 5) 1 $>$ greatest_common_divisor(25, 15) 5 """
@@
Table 4: Task prompt for per-task pruning experiments.
A.4 Submission Checklist

### v3

@@ 4.2 Evaluation with Input-Specific Masks
Second, IFPruning demonstrates superior performance compared to the structured pruning method. Please note that the structured pruning model, Pruning+Distill, benefits from additional training signals due to knowledge distillation from a larger teacher model as the teacher. Nevertheless, IFPruning consistently outperforms this baseline. Our method achieves higher performance across a variety of benchmarks, including AlpacaEval, Arena Hard, math problems, coding tasks, tool use, MMLU, and most Core Text tasks.
Third, we observe a clear improvement in performance with IFPruning as the size of the LLM increases. As the source model size scales from 6B to 9B and then to 12B, there is a noticeable performance boost on most of the datasets.
@@
Table 2: Performance comparison between IFPruning with per-input masks and per-task masks. In the per-input setting, IFPruning prunes the LLMs and activates different sub-networks for each input. In the per-task setting, IFPruning selects a single sub-network for all inputs within the same dataset based on a human-written task instruction.
@@
We highlight the following findings. First, in the lower layers, especially the first layer, the LLM tends to activate very similar sub-networks for different inputs. As we move to higher layers, the parameter selection becomes more diverse. Second, as shown in the figure, IFPruning activates distinct sub-networks for different domains in higher layers. For instance, models pruned for MMLU-CS have substantial overlap with those for Code-Alpaca, significantly more than with other domains. Similarly, models for MMLU-Math, MMLU-Physics, and GSM8K share a high proportion of activated parameters, which diverge notably from the activation patterns for MMLU-History, MMLU-Law, and GPTTeacher. Third, the overlap rate along the diagonal of the heatmap is very high, reflecting the strong similarity between models for inputs within the same domain. Finally, as expected, GPTTeacher is an instruction-following dataset that covers a wide range of domains, therefore it does not exhibit significant domain-specific characteristics, resulting in a self-overlap rate that is lower than in other domains. In summary, the activated parameters are interpretable and reveal clear patterns in different domains. It aligns with the high performance of our method, as the LLM dynamically activates the parameters most suitable for each input.
Ablation study
We also conduct an ablation study on the impact of continued pre-training in Appendix A.4. Incorporating continued pre-training before SFT leads to consistent and substantial improvements across all benchmarks.
@@ 4.3 Evaluation with Task-Specific Masks
An additional noteworthy capability of IFPruning is task-specific pruning. While IFPruning can effectively prune a model on a per-input basis, the algorithm also demonstrates the ability to select one sub-network that can be used for different inputs within the same task or domain. We design the following experiment to demonstrate the performance of our method with task-specific pruning.
Specifically, we evaluate our models on math problems, coding tasks, MMLU, and machine translation tasks from Flores-101 (Goyal et al., 2022). For each dataset, we manually design a concise instruction that describes the task. For example, the instruction for the HumanEval dataset can be: “You are an expert Python programmer. Write the code which should pass the tests.” The sparsity predictor takes this instruction as input and selects the corresponding sub-network. This sub-network is then applied consistently to process all testing examples from that dataset, and the model performance on the dataset is evaluated based on the predictions made by the sub-network. The task-specific instructions used for each dataset are provided in Table 6 in Appendix A.7. For the MMLU dataset, which consist of multiple subsets representing diverse domains, we group similar subsets into broader domains and assign a single instruction to each domain. For example, the subsets “astronomy,” “college physics,” “conceptual physics,” and “high school physics” are grouped into the domain “MMLU-physics,” and a single instruction is written for this domain. A detailed list of the subsets included in each MMLU domain is provided in Appendix A.6.
@@
The performance is shown in Table 2. We compare the task-specific pruning capabilities of our algorithm with the dense LLM with 3B parameters and the standard IFPruning with per-input mask. Our key findings are as follows. First, IFPruning can generate high-quality task-specific masks without additional training. we observe substantial performance improvements of the task-specific IFPruning over the dense baseline across all datasets. When applying per-task masks, IFPruning still outperforms the dense LLM by 5%-12% on coding tasks, 5%-6% on MATH, and 1% - 5% on MMLU subsets. For translation tasks, IFPruning selects the appropriate sub-networks without additional training, achieving an improvement of 4.8 points in BLEU score on average. Note that these task-specific masks are directly predicted by the sparsity predictor, requiring no additional fine-tuning of either the sparsity predictor or the LLM for each task. This highlights the “zero-shot” pruning capability of our method. Second, compared to the standard input-specific IFPruning, the task-specific IFPruning exhibit minimal performance degradation across math problems, coding tasks, and MMLU. The performance gaps are mostly under 1%, indicating the robustness of our method.
4.4 Inference Latency Evaluation
Given the strong performance of IFPruning, we further analyze and evaluate its efficiency improvement in this section. The inference pipeline of our method consists of three steps: ❶ Parameter selection: given an input, the sparsity predictor selects a subset of parameters to activate. ❷ Parameter loading: the selected parameters are loaded, and the corresponding sub-network is cached as a dense model. ❸ Generation: the pruned model will encode the input (prefill) and generate a response (decode), following the same procedure as standard dense or MoE models. Sub-network selection only involves a forward pass through the sparsity predictor, which is lightweight (300M parameters in our experiments), so the time cost is minimal. For parameter loading, we maintain a small base model and update its weights with the selected parameters. The generation step is identical to standard inference, allowing our method to achieve the same latency as a smaller dense model.
[…]
We highlight the following conclusions. First, similar to standard structured pruning, our dynamic pruning scheme introduced can also significantly reduces inference latency for large-scale models. By pruning a 9B or 8B LLM to 3B, the TTFT is decreased by up to 57% and decoding time is decreased by up to 41%. In total, IFPruning can achieve up to 1.8x speedup compared to the original dense LLMs. Second, the overhead introduced by dynamic pruning and parameter caching is negligible. For each input, the parameter selection and loading only take less than 0.1s in total (1-2% of the total generation time). Finally, despite dynamic masking, the runtime of IFPruning is on par with static pruning (Dense-3B baseline), while offering input-specific adaptivity and superior accuracy. The inference latency evaluation has further verified that our method is suitable for practical deployment on edge devices, as it accelerates inference while achieving superior model performance.
Finally, we compare the inference latency of our method and an MoE model with similar number of activated parameters in Appendix A.5. We show that for the on-device inference scenario with small inference batch size, the cost of MoE is multiple times higher than a dense model and our method.
Table 3: Inference latency evaluation. The source models are reduced to 3B by pruning the FFN layers.
Model Parameter Selection Parameter Loading Prefill TTFT Decode Our model Dense-9B - - 0.846 0.846 7.16 Dense-3B - - 0.356 0.356 5.68 IFPruning 0.070 0.043 0.358 0.471 5.58 Llama-3.1-8b Dense-8B - - 0.702 0.702 5.47 Dense-3B - - 0.317 0.317 3.52 IFPruning 0.070 0.016 0.315 0.402 3.53
@@ 5 Conclusion
This work also opens several promising directions for future exploration. First, we focus on dynamically pruning LLMs based on contextual information. While this approach is well suited for models on consumer-facing devices such as phones, additional challenges remain for server-side serving when the input is a batch of user requests containing different tasks. One possible solution is to cluster user requests so that requests within the same batch share similar activated sub-networks. Second, the current training method relies on end-to-end optimization, which may not fully utilize the training examples. The performance could be improved by adopting more advanced training strategies. For instance, using contrastive loss could encourage higher overlap rates among sub-networks for similar inputs, making the model more robust.
Acknowledgements
Shiyu Chang acknowledges support from National Science Foundation (NSF) Grant IIS-2338252, NSF Grant IIS-2207052, and NSF Grant IIS-2302730.
Impact Statement
@@ Acknowledgements
Appendix A Appendix
A.1 Implementation Details
LLM architecture
The LLMs used in this paper follow standard LLM design (Meta AI, 2024; Team, 2024), such as grouped-query attention and RMSNorm, with no custom components. We list the detailed model architecture of the pre-trained models used in our experiments in Table 4. All models use the same model dimension, attention dimensions and the number of Transformer layers. The only difference is the feed-forward dimension. Accordingly, our method learns to select 6656 FFN dimensions from the 6B and 9B model.
@@ A.1 Implementation Details
Sparsity predictor
The predictor is built upon a lightweight LLM with 302M parameters, built on a pre-trained LM backbone. It consists of:
  • A small pre-trained language model with 302M parameters as the feature extractor. Given an input, the last hidden state of the final input token will be used as the features of the input.
  • A Two-layer MLP as the prediction head. Given the extracted feature vector with dimension $d$, the first layer transform the input from $d$ to 128. The second layer transform the output of the first layer into $L*d_{\mathrm{ffn}}$, where $L$ is the number of layers of the source LLM being pruned and $d_{\mathrm{ffn}}$ is the intermediate dimension of the source LLM. Then we reshape the output vector into a $l$-by-$d_{\mathrm{ffn}}$ matrix. Each row of the matrix will be processed by the SoftTopK (Lei et al., 2023) operator independently to get the FFN mask for the corresponding layer.
A.2 Training of the Structured Pruning Baseline
Our pruning baseline, Pruning+Distill, first prunes the given LLM into a dense 3B model then perform continuous pre-training through knowledge distillation. For knowledge distillation. we apply KL divergence between the output distributions of the student and the teacher model for each output token. The teacher distribution only keeps the highest-scoring tokens, similar to Minitron (Muralidharan et al., 2024). We minimize a combined loss of the standard next-token prediction and KL divergence loss.
A.3 Interpretability of Parameter Selection
We provide more details of the visualization of the parameter selection pattern in Figure 3. First, we sample 128 inputs per dataset (MMLU, GSM8K, CodeAlpaca-20K, GPTTeacher). Inputs to the sparsity predictor are formatted with in-context examples (MMLU: 5-shot, GSM8K: 8-shot) or raw prompts (CodeAlpaca, GPTTeacher). After that, each input is passed through the sparsity predictor, which selects a fixed number of FFN units (a binary mask) at each layer, producing 128 sub-networks per dataset. To calculate the overlap, We compare every pair of sub-networks within the 128 examples. For each pair, we compute the fraction of selected FFN units they share. The final overlap rate is the average of these pairwise overlaps.
A.4 Ablation Study on the Continued Pre-training
In this section, we conduct a ablation study on the impact of the continued pre-training. We apply our method and activate 3B parameters of a Dense-6B LLM. We train two variants of models: one with the SFT phase ony and one with both continued pre-training and SFT phases. The experiment results are shown in Table 5. We observe consistent and notable gains across all benchmarks, demonstrating the effectiveness of the continued pre-training phase.
Table 5: Ablation study on the impact of the continued pre-training. We apply IFPruning to a dense LLM with 6B parameters and activate 3B parameters for each input.
HumanEval MBPP MultiPL-E GSM8K MATH MMLU
SFT Only 25.3 24.4 15.3 50.9 13.5 55.2
Continued pre-training + SFT 31.9 35.3 22.4 61.3 20.1 59.3
A.5 Inference Latency Comparison with MoE
Figure 4: Inference latency evaluation. The source models are reduced to 3B by pruning the FFN layers.
Model Param. selection Param. loading Prefill TTFT Decode Dense-9B - - 0.846 0.846 7.16 MoE-A2.7B - - 0.621 0.621 28.43 Dense-3B - - 0.356 0.356 5.68 IFPruning 0.070 0.043 0.358 0.471 5.58
In this section, we compare the latency of our method with the open-source model, Qwen1.5-MoE-A2.7B. It has 14.3B parameters in total and activates 2.7B parameters per token. For our method, we prune LLaMA-3-8B to 3B parameters. The experiment configurations are exactly the same as the experiment in Table 3. Specifically, we evaluate the latency on a single NVIDIA RTX A6000 GPU and report time-to-first-token (TTFT) and decoding time with input length = 4k, generation length = 100, and sample 4 responses for each query. The experiment results are shown in Table 4.
We can see the dense baseline and our method have significantly better latency and throughput than MoE. The main reason is that MoE models are not efficient for on-device inference when generating responses given a single query, where the decoding is bottlenecked by weight loading. Since MoE requires reading many expert weights (e.g., 2 for each token), the cost of MoE is multiple times higher than a dense model and our method. This further demonstrates that our method is more suited for on-device scenarios.
A.6 Subsets included in MMLU domains
@@
A.7 Detailed Per-task Prompts
@@
Table 6: Task prompt for task-specific pruning experiments.
@@
Multiple-E Swift You are an expert Swift programmer. You will be given a Swift function definition in documentation comments after ///. Write the code to complete the function. Here is an example input: ```swift /// Write a swift function to count inversions in an array. func get_Inv_Count(arr: [Int]) $\rightarrow$ Int
HumanEval-Python You are an expert Python programmer. You will be given a Python function definition and some test examples in triple quotes """. Write the code which should pass the tests. Here is an example input: ```python def greatest_common_divisor(a: int, b: int) -¿ int: """Return a greatest common divisor of two integers a and b $>$ greatest_common_divisor(3, 5) 1 $>$ greatest_common_divisor(25, 15) 5 """
@@
A.8 License of Datasets and Use of Generative AI